# Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness

### v1

@@ C.2 Qualitative Analysis
We provide comprehensive quantitative evaluations including additional sensitivity analysis across hyperparameters.
@@ Appendix D Statement on the Use of Large Language Models
Large Language Models (LLMs) were used in limited capacity during the preparation of this research. Specifically, LLMs were used to check grammar and refine sentence structure after the initial draft was completed, primarily to correct awkward expressions and maintain consistency in writing style. However, all core research ideas, analytical methodologies, interpretations of the results, and conclusions were developed entirely by the authors. The LLM did not contribute to any creative content or academic judgments. This use of LLMs was conducted within limits that do not compromise the originality or academic integrity of the research.

### v2

@@ C.2 Qualitative Analysis
We provide comprehensive quantitative evaluations, including additional sensitivity analysis across hyperparameters.
@@ Appendix D Statement on the Use of Large Language Models
Large Language Models (LLMs) were used in a limited capacity during the preparation of this research. Specifically, LLMs were used to check grammar and refine sentence structure after the initial draft was completed, primarily to correct awkward expressions and maintain consistency in writing style. However, all core research ideas, analytical methodologies, interpretations of the results, and conclusions were developed entirely by the authors. The LLM did not contribute to any creative content or academic judgments. This use of LLMs was conducted within limits that do not compromise the originality or academic integrity of the research.